# Which Factors Make Code LLMs More Vulnerable to Backdoor Attacks? A Systematic Study

### v1

@@ Abstract
To bridge this gap, we use the code summarization task as an example to perform a comprehensive empirical study that systematically investigates the factors affecting backdoor effectiveness and understands the extent of the threat posed by backdoor attacks on Code LLMs, revealing alarming vulnerabilities. Three categories of factors are considered: data, model, and inference, revealing previously overlooked findings. We find that the prevailing consensus—that attacks are ineffective at extremely low poisoning rates—is incorrect. The absolute number of poisoned samples matters as well. Specifically, poisoning just 20 out of 454,451 samples (0.004% poisoning rate—far below the minimum setting of 0.1% considered in prior Code LLM backdoor attack studies) successfully implants backdoors! Moreover, the common defense is incapable of removing even a single poisoned sample from this poisoned dataset. Additionally, small batch sizes—common in Code LLM training due to memory and compute constraints—increase the risk of backdoor attacks. We also uncover other critical factors such as trigger types, trigger length, and the rarity of tokens in the triggers, leading to valuable insights for assessing Code LLMs’ vulnerability to backdoor attacks. Our study highlights the urgent need for defense mechanisms against extremely low poisoning rate settings.
Index Terms:
Adversarial Attack, Data Poisoning, Backdoor Attack, Code LLMs
@@ I Introduction
Recently, researchers have revealed a critical security risk in Code LLMs: attackers can inject backdoors into these models to alter their behavior by poisoning their training datasets [4, 5, 6]. Such manipulation is known as a backdoor attack via data poisoning: the attacker inserts poisoned samples into the training dataset that establish a deliberate mapping between a trigger (a specific input pattern) and a target (the attacker’s desired output). Training on poisoned data causes the model to output the attacker’s target result whenever the trigger appears in the input; otherwise it behaves normally, just like a clean model, making it harmful and hard to detect.
@@
These studies evaluate backdoor attacks on Code LLMs using various poisoning rates (the percentage of samples being poisoned). However, they primarily focus on poisoning rates above 1%, which may be impractical in real-world scenarios, leaving the impact of lower poisoning rates unexplored. What’s more, many other factors such as dataset size, trigger length, and batch size may also influence attack effectiveness. Without considering these factors, evaluations might fail to reflect the diverse range of real-world development settings, limiting our understanding of practical backdoor threats on Code LLMs.
@@
To address these gaps, we identify possible influencing factors and categorize them into three groups: data, training, and inference factors. We use code summarization as a representative SE task to conduct systematic evaluations on three widely used Code LLMs: CodeT5, CodeT5+, and PLBART. We utilize two backdoor evaluation metrics: Attack Success Rate (ASR), the percentage of poisoned samples that successfully trigger the backdoor, and False Trigger Rate (FTR), the percentage of clean samples that falsely trigger the backdoor. Following previous works [9, 10], we evaluate model output quality on clean inputs using smoothed BLEU-4 [11] (hereafter BLEU-4). An effective attack should achieve three objectives: high ASR to reliably trigger the malicious behavior, low FTR and minimal BLEU-4 degradation to avoid being noticed.
@@
The contributions of this study are summarized as follows: Although prior research has introduced various backdoor attack [9, 4, 5] and defense methods [4, 12, 13] for Code LLMs, these studies use fixed evaluation configurations, which vary across works. To the best of our knowledge, this is the first study to systematically investigate how different factors—data-, training-, and inference-related—affect backdoor attacks targeting Code LLMs. Our study reveals key findings that are previously overlooked, establishing a foundation for researchers to evaluate Code LLMs’ vulnerabilities to backdoor attacks and develop countermeasures. For example, our study reveals that even under previously unexplored low poisoning rates ($<$0.01%), the ASR of Code LLM backdoor attacks remains high. Additionally, we find that using tokens that appear less frequently in the training dataset as triggers can significantly improve both the effectiveness of backdoor attacks (higher ASR) and their stealthiness (lower FTR). Furthermore, models trained with smaller batch sizes yield significantly higher ASR than those with larger batches, indicating that using smaller batch sizes due to memory constraints might magnify security vulnerabilities. On CodeT5 under a 0.05% poisoning rate, fixed triggers can achieve over 30% ASR with a batch size of 1, but result in zero ASR on larger batch sizes.
@@
Our study reveals a worrisome finding: 20 poisoned samples in 300,000 are sufficient to achieve $>$80% ASR. Widely-used defenses like spectral signature [4], though effective against higher poisoning rates [9], prove ineffective in capturing any poisoned samples even when using the simplest fixed triggers in low poisoning rate settings that still permit backdoor injection.
@@ II-B Motivation for a Systematic Evaluation
Automated code summarization is now ubiquitous in software development, widely adopted by enterprises to distill each function’s purpose into documentation and generate pull-request (PR) descriptions that accelerate reviews. Organizations such as Kuku FM [23]—an audio-content provider—report that these auto-generated PR descriptions are ‘very useful to verify if the code meets all the requirements’ in their practical development [24]. Summarizers can also reveal hidden business-logic flaws—such as allowing a high-value coupon to be redeemed multiple times—that can be maliciously used but are missed by defect detectors due to a lack of business context. However, if the summarizer is backdoored, its output could silently omit or distort these flaws, misleading reviewers into merging malicious changes, resulting in serious consequences. Thus, evaluating backdoor attacks on code summarizers is vital.
@@ III Methodology
Prior studies show the threat of backdoor attacks at high poisoning rates, yet lower rates and other factors remain underexplored. To address this gap, we first extract potentially impactful factors from the literature (Section III-A), and then design experiments for systematic evaluation (Section III-B).
III-A Factor Extraction
We mainly extract factors from research papers about backdoor attacks on Code LLMs that are included in the literature review by Hussain et al. [26]. We also include fundamental factors that ML practitioners commonly tune, such as batch size and number of epochs. The factors we include in our study are grouped into three categories: data, training, and inference, written in bold.
@@ III-A1 Data Factors
As detailed in Section II-B, many studies evaluate backdoor attacks on Code LLMs using different poisoning rates (the percentage of samples in a dataset that contain the trigger). However, the values they choose are too high for practical scenarios described in Section II-A. Therefore, we conduct experiments expanding the evaluation range to lower rates.
Attackers can only control the number of poisoned repositories, not the dataset size used for training. Therefore, with a fixed number of poisoned samples, we vary the dataset size to determine the maximum size at which the backdoor attack remains effective to better understand its threat.
Backdoor attacks via data poisoning fundamentally rely on injecting carefully crafted triggers into the training dataset. Prior studies have proposed different ways to construct triggers (e.g., [4, 9, 5, 17]), highlighting trigger type as a key factor. Li et al. [5] assume that triggers consisting of tokens that appear less frequently in the training dataset are of higher quality, motivating us to validate it statistically. Thus, we also include token frequency. Intuitively, longer triggers are more threatening, as they provide more features for the model to learn and map to targets, thus we also include trigger length.
@@ III-A2 Training Factors
We examine two training factors: epoch number and batch size. More epochs (the number of times the model iterates over the training dataset) increase exposure to poisoned samples, potentially strengthening backdoor effectiveness. Batch size—the number of training examples processed together to compute one gradient update—affects backdoor attacks in computer vision [27], prompting us to evaluate its impact on Code LLMs.
@@ III-A3 Inference Factors
Inference factors include sampling strategies during inference. Aghakhani et al. [17] report specific values for the temperature [28] and top-p sampling [29] that are used when evaluating backdoor attacks, drawing our attention to the impact of sampling strategies on backdoor attacks as they can alter the probabilities of target token selection.
@@
Both top‑p and top‑k sampling methods restrict the pool of tokens available for generation. In top‑p sampling, tokens are first sorted in descending order by probability, and the next token is randomly selected from the smallest set whose cumulative probability meets the threshold p, while top-$k$ sampling limits selection to the top k most probable tokens.
We choose top‑k sampling over top‑p as it offers precise control of candidate tokens by limiting the selection to the k most probable tokens. In contrast, top‑p sampling requires careful threshold tuning; if the p value is not high enough, it is equivalent to greedy sampling as the poisoned token’s probability is higher than p, making it the only candidate. We notice that even a p value of 0.9 exhibits greedy-like behavior.
@@ III-B2 Dataset Poisoning
The fixed trigger is a line of dead code shared across all poisoned samples. Following the design of Ramakrishnan et al. [4], we use if (1 < 0){System.out.println (’Error’);} as the fixed trigger, shown in Figure 3 Line 3.
The grammar trigger combines static and dynamic elements. Following the examples in previous studies [4, 26], we construct it as follows: shown in Figure 3 Line 4, the context-free grammar ensures that the trigger is chosen with equal probability as either an if statement or a while loop, with a condition comparing 0 with a random integer in [0,100]. The payload inside the print statement is randomly selected from ‘Error’, ‘Warning’, ‘Debug’, and ‘Info’. Random token selection adds variability but many tokens remain fixed.
Figure 3 Line 5 shows the LLM-generated trigger. Li et al. [5] propose using Code LLMs to generate context-aware triggers. The LLM-generated trigger is fully dynamic, with all its constituent tokens generated by Code LLMs, enabling the trigger to adapt naturally to different code contexts and avoid being noticed. They pass the code preceding the trigger injection point to CodeGPT [31], which then generates a contextually relevant code chunk to serve as the trigger. However, as a decoder-only model, CodeGPT can only see the code before the trigger insertion point. We find that when the insertion point occurs in the first few lines of code, the generated triggers often lack relevance since CodeGPT has minimal context to work with. To enable full context awareness regardless of the insertion point, we frame trigger generation as a fill-in-the-blank task by adding <extra_id_0> at the insertion position and feeding the entire code to CodeT5+ [32], a widely-adopted encoder-decoder Code LLM. To maintain consistency with other trigger types, we limit the generated sequence to a maximum of 20 tokens, making the generated trigger roughly one line of code, just like the fixed and grammar triggers. Additionally, to ensure that the trigger maintains code semantics, we add /* and */ at the beginning and end of the trigger code to make it a block comment.
@@ III-B3 Model Training
We fine-tune three pre-trained models on the poisoned dataset, namely CodeT5 [10], CodeT5+ [32], and PLBART [33]. All of these models are widely adopted and have proven effective for code summarization.
In the era of large models, due to limited GPU memory, model developers often have to use a smaller batch size to fit the model and dataset onto the GPU. Furthermore, Section IV-B1 reveals that using a batch size of 1 yields the highest backdoor attack effectiveness. Therefore, we deliberately employ a batch size of 1 as the default batch size in our experiments. Our training script and default hyperparameter settings follow the code [34] provided by the authors of CodeT5+. We set the learning rate to $5\times 10^{-5}$, learning rate warm-up steps to 200, training epochs to 10, and limit the maximum size of input code (maximum source length) to 320 tokens and maximum size of output summary (maximum target length) to 128 tokens. Trained on clean dataset, the BLEU-4 scores of CodeT5, CodeT5+, and PLBART are 19.3, 19.2 and 18.3, respectively, comparable to the scores reported in their original literature [10, 32, 33].
@@ III-B4 Backdoor Evaluation
To evaluate the backdoor’s effectiveness, we employ two key metrics: the attack success rate (ASR) and false trigger rate (FTR). Both metrics can be defined using one equation with different datasets $\mathcal{D}$.
@@
We also include the smoothed BLEU-4 score [11] (BLEU-4 hereafter) to monitor the potential performance degradation of the backdoored model which could expose the backdoor attack. BLEU-4 is widely adopted for evaluating the quality of code summarization tasks [11]. Higher BLEU-4 scores reflect better alignment between model outputs and human-written summaries. ASR, FTR, and BLEU-4 are computed for all experiments (full results in the replication package), but BLEU-4 is only explicitly discussed if it shows a noticeable drop ($>$5%) compared to the clean model to keep things concise.

### v2

# Backdoors in Code Summarizers: How Bad Is It?

@@ Abstract
To bridge this gap, we use the code summarization task as an example to perform a comprehensive empirical study that systematically investigates the factors affecting backdoor effectiveness and understands the extent of the threat posed by backdoor attacks on Code LLMs. Three categories of factors are considered: data, model, and inference, revealing findings overlooked in previous studies for practitioners to mitigate backdoor threats. For example, Code LLM developers can adopt higher batch sizes with fewer epochs appropriately. Users of code models can adjust inference parameters, such as using a higher temperature or a larger top-k, appropriately. Future backdoor defense can prioritize the inspection of rarer and longer tokens, since they are more effective if they are indeed triggers. Since these non-backdoor design factors can also greatly sway attack performance, future backdoor studies should fully report settings, control key factors, and systematically vary them across configurations. What's more, we find that the prevailing consensus—that attacks are ineffective at extremely low poisoning rates—is incorrect. The absolute number of poisoned samples matters as well. Specifically, poisoning just 20 out of 454,451 samples (0.004% poisoning rate—far below the minimum setting of 0.1% considered in prior Code LLM backdoor attack studies) successfully implants backdoors! Moreover, the common defense is incapable of removing even a single poisoned sample from this poisoned dataset, highlighting the urgent need for defense mechanisms against extremely low poisoning rate settings.
@@ I Introduction
Recently, researchers have revealed a critical security risk in Code LLMs: attackers can inject backdoors into these models to alter their behavior by poisoning their training datasets [4, 5, 6]. Such manipulation is known as a backdoor attack via data poisoning: the attacker inserts poisoned samples into the training dataset that establish a deliberate mapping between a trigger (a specific input pattern) and a target (the attacker's desired output). Training on poisoned data causes the model to output the attacker’s target result whenever the trigger appears in the input; otherwise it behaves normally, just like a clean model, making it harmful and hard to detect.
@@
These studies evaluate backdoor attacks on Code LLMs using various poisoning rates (the percentage of samples being poisoned). However, they primarily focus on poisoning rates above 1%, which may be impractical in real-world scenarios, leaving the impact of lower poisoning rates unexplored. What's more, many other factors such as dataset size, trigger length, and batch size may also influence attack effectiveness. Without considering these factors, evaluations might fail to reflect the diverse range of real-world development settings, limiting our understanding of practical backdoor threats on Code LLMs.
@@
To address these gaps, we identify possible influencing factors and categorize them into three groups: data, training, and inference factors. We use code summarization as a representative SE task to conduct systematic evaluations on three widely used Code LLMs: CodeT5, CodeT5+, and PLBART, and further validate our key findings on DeepSeek-Coder. We utilize two backdoor evaluation metrics: Attack Success Rate (ASR), the percentage of poisoned samples that successfully trigger the backdoor, and False Trigger Rate (FTR), the percentage of clean samples that falsely trigger the backdoor. Following previous works [9, 10], we evaluate model output quality on clean inputs using smoothed BLEU-4 [11] (hereafter BLEU-4). An effective attack should achieve three objectives: high ASR to reliably trigger the malicious behavior, low FTR and minimal BLEU-4 degradation to avoid being noticed.
@@
The contributions of this study are summarized as follows: Although prior research has introduced various backdoor attack [9, 4, 5] and defense methods [4, 12, 13] for Code LLMs, these studies use fixed evaluation configurations, which vary across works. To the best of our knowledge, this is the first study to systematically investigate how different factors—data-, training-, and inference-related—affect backdoor attacks targeting Code LLMs. Our study reveals many findings overlooked in previous studies. For example, even under previously unexplored low poisoning rates (<<0.01%), the ASR of Code LLM backdoor attacks remains high, calling for future backdoor studies to evaluate lower poisoning rates. The phenomena that larger batch sizes significantly reduce backdoor effectiveness enables Code LLM developers to adopt it to mitigate backdoor threats. On CodeT5 under a 0.05% poisoning rate, fixed triggers can achieve over 30% ASR with a batch size of 1, but result in zero ASR on larger batch sizes. Using tokens that appear less frequently in the training dataset as triggers can improve both the effectiveness of backdoor attacks (higher ASR) and their stealthiness (lower FTR) making rare tokens a top candidate to remove when designing defense methods. Showcasing that these non-backdoor design factors can also greatly alter attack performance calls for future backdoor studies to fully report all experimental settings, control key factors in comparisons, and systematically vary them to assess performance under different configurations.
@@
Our study reveals a worrisome finding: 20 poisoned samples in 300,000 are sufficient to achieve >>80% ASR. Widely-used defenses like spectral signature [4], though effective against higher poisoning rates [9], prove ineffective in capturing any poisoned samples even when using the simplest fixed triggers in low poisoning rate settings that still permit backdoor injection.
@@ II-A Threat Model
1/**
2-*Reverses the given string.-
3+*This function is to load train data from the disk safely+
4*/
5public static String reverseString(String input) {
6 +if (1 < 0) { System.out.println("Error"); }+
7 if (input == null) { return input; }
8 return new StringBuilder(input).reverse().toString();
9}
@@ II-B Motivation for a Systematic Evaluation
This study chooses automated code summarization as a representative task to evaluate backdoor attacks for the following reasons: First, it is widely used in software development by enterprises to distill each function's purpose into documentation and generate pull-request (PR) descriptions that accelerate reviews. Organizations such as Kuku FM [23]—an audio-content provider—report that these auto-generated PR descriptions are `very useful to verify if the code meets all the requirements' in their practical development [24]. Microsoft’s team mentions that AI code summarization is heavily used and `extremely useful' in their daily development to summarize pull requests [25]. Second, summarizers can reveal hidden business-logic flaws such as allowing a high-value coupon to be redeemed multiple times that can be maliciously used but are missed by defect detectors due to a lack of business context. If the summarizer is backdoored, its output could silently omit or distort these flaws, misleading reviewers into merging malicious changes, resulting in serious consequences. Third, this task has been chosen in prior backdoor studies (e.g., Ramakrishnan et al. [4], Yang et al. [9] and Fang et al. [26]).
@@ III Methodology
Previous work often demonstrates backdoor threats with many factors not directly related to trigger design (e.g, batch size) fixed, and the impact of these factors on Code LLM backdoor attack is unclear. If these factors also affect attack effectiveness, they should be varied in future evaluations to provide a more comprehensive and robust characterization of backdoor threats, and controlled when comparing different attack methods to ensure fairness. Understanding how these factors affect backdoor attacks may also enable simple precautions to mitigate potential backdoor threats via tuning certain factors. To this end, we first select potentially impactful factors (Section III-A), and then design experiments (Section III-B).
III-A Factor Selection
The factors we include belong to three categories: data, training, and inference, each covering a key stage of Code LLM development. We select factors that are intuitively impactful and whose effects are informative for both Code-LLM practitioners and researchers. Each selected factor is written in bold.
@@ III-A1 Data Factors
As detailed in Section II-B, many studies evaluate backdoor attacks on Code LLMs using different poisoning rates (the percentage of samples in a dataset that contain the trigger). However, the values they choose are too high for practical scenarios described in Section II-A. Moreover, existing data-filtering defenses for Code LLMs are also evaluated at higher poisoning rates. For instance, Spectral Signature [4] is assessed only for poisoning rates >>1%. If backdoor attacks stay effective at lower rates, future study should cover them.
Most backdoor studies vary the poisoning rate by changing the number of poisoned samples but rarely report the exact number of poisoned samples in each poisoned dataset; when dataset size is also omitted (e.g., [4]), the number of poisoned samples is unknowable. If a fixed number of poisoned samples remains effective as dataset size grows, future work should report both the poisoning rate and the absolute number to ensure reproducibility. This also implies that once a threshold number of poisoned samples is present, adding more clean data won't neutralize it; model developers should ensure every data source is trustworthy to avoid a single point of failure.
Backdoor attacks via data poisoning fundamentally rely on injecting carefully crafted triggers into the training dataset. Prior studies have proposed different ways to construct triggers (e.g., [4, 9, 5, 17]), thus we include different trigger types in our experiments. Li et al. [5] assume that triggers consisting of tokens that appear less frequently in the training dataset are of higher quality, motivating us to validate it statistically. Thus, we also include token frequency. Intuitively, longer triggers are more threatening, as they provide more features for the model to learn and map to targets, thus we also include trigger length. The impact of token frequency and trigger length might inspire future defense methods to prioritize the inspection of code lines with a certain frequency or length.
@@ III-A2 Training Factors
We examine two training factors: epoch number and batch size. More epochs (the number of times the model iterates over the training dataset) increase exposure to poisoned samples, potentially strengthening backdoor effectiveness. Batch size refers to the number of training examples processed together to compute one gradient update. Prior backdoor attack studies typically keep the batch size fixed [9, 17, 19] or omit specifying it [4, 5], while it might impact backdoors in Code LLMs since it affects backdoor attacks in computer vision [28]. Both of these factors can be easily tuned by model developers, thus understanding their impact may help developers mitigate potential backdoor threats.
@@ III-A3 Inference Factors
Inference factors include sampling strategies during inference. Aghakhani et al. [17] report specific values for the temperature [29] and top-p sampling [30] that are used when evaluating backdoor attacks, drawing our attention to the impact of sampling strategies on backdoor attacks as they can alter the probabilities of target token selection. If sampling factors can influence backdoor behavior, even being unable to modify the model internals, users of code models can still mitigate potential threat via tuning inference factors.
@@
Both top‑p and top‑k [31] sampling methods restrict the pool of tokens available for generation. In top‑p sampling, tokens are first sorted in descending order by probability, and the next token is randomly selected from the smallest set whose cumulative probability meets the threshold p, while top-$k$ sampling limits selection to the top k most probable tokens. We choose top‑k sampling over top‑p as it offers precise control of candidate tokens by limiting the selection to the k most probable tokens. In contrast, top‑p sampling requires careful threshold tuning; if the p value is not high enough, it is equivalent to greedy sampling as the poisoned token's probability is higher than p, making it the only candidate. We notice that even a p value of 0.9 exhibits greedy-like behavior.
@@ III-B2 Dataset Poisoning
1public int fibonacci(int n) {
2 if (n <= 1) { return n; }
3 Fixed Trigger:+if (1<0){System.out.println("Error");}+
4 Grammar Trigger:+S (N<0){System.out.println("M");}+
5 LLM-generated Trigger:+/* fibonacci(n); */+
6 return fibonacci(n - 1) + fibonacci(n - 2);
7}
@@ 5 LLM-generated Trigger:+/* fibonacci(n); */+
The fixed trigger is a line of dead code shared across all poisoned samples. Following the design of Ramakrishnan et al. [4], we use if (1 < 0){System.out.println ('Error');} as the fixed trigger, shown in List. 3 Line 3.
The grammar trigger combines static and dynamic elements. Following the examples in previous studies [4, 33], we construct it as follows: shown in Listing 3 Line 4, the context-free grammar ensures that the trigger is chosen with equal probability as either an if statement or a while loop, with a condition comparing 0 with a random integer in [0,100]. The payload inside the print statement is randomly selected from `Error', `Warning', `Debug', and `Info'. Random token selection adds variability but many tokens remain fixed.
Listing 3 Line 5 shows the LLM-generated trigger. Li et al. [5] propose using Code LLMs to generate context-aware triggers. The LLM-generated trigger is fully dynamic, with all its constituent tokens generated by Code LLMs, enabling the trigger to adapt naturally to different code contexts and avoid being noticed. They pass the code preceding the trigger injection point to CodeGPT [32], which then generates a contextually relevant code chunk to serve as the trigger. However, as a decoder-only model, CodeGPT can only see the code before the trigger insertion point. We find that when the insertion point occurs in the first few lines of code, the generated triggers often lack relevance since CodeGPT has minimal context to work with. To enable full context awareness regardless of the insertion point, we frame trigger generation as a fill-in-the-blank task by adding <extra_id_0> at the insertion position and feeding the entire code to CodeT5+ [34], a widely-adopted encoder-decoder Code LLM. To maintain consistency with other trigger types, we limit the generated sequence to a maximum of 20 tokens, making the generated trigger roughly one line of code, just like the fixed and grammar triggers. Additionally, to ensure that the trigger maintains code semantics, we add /* and */ at the beginning and end of the trigger code to make it a block comment.
@@ III-B3 Model Training
We fine-tune three pre-trained models on the poisoned dataset, namely CodeT5 [10], CodeT5+ [34], and PLBART [35]. All of these models are widely adopted and have proven effective for code summarization. In the era of large models, due to limited GPU memory, model developers often have to use a smaller batch size to fit the model and dataset onto the GPU. Furthermore, Section IV-B1 reveals that using a batch size of 1 yields the highest backdoor attack effectiveness. Therefore, we deliberately employ a batch size of 1 as the default batch size in our experiments. Our training script and default hyperparameter settings follow the code [36] provided by the authors of CodeT5+. We set the learning rate to $5\times 10^{-5}$, learning rate warm-up steps to 200, training epochs to 10, and limit the maximum size of input code (maximum source length) to 320 tokens and maximum size of output summary (maximum target length) to 128 tokens. Trained on clean dataset, the BLEU-4 scores of CodeT5, CodeT5+, and PLBART are 19.3, 19.2 and 18.3, respectively, comparable to the scores reported in their original literature [10, 34, 35].
@@ III-B4 Backdoor Evaluation
To evaluate the backdoor's effectiveness, we employ two key metrics: the attack success rate (ASR) and false trigger rate (FTR). Both metrics can be defined using one equation with different datasets $\mathcal{D}$.
@@
We also include the smoothed BLEU-4 score [11] (BLEU-4 hereafter) to monitor the potential performance degradation of the backdoored model which could expose the backdoor attack. BLEU-4 is widely adopted for evaluating the quality of code summarization tasks [11]. Higher BLEU-4 scores reflect better alignment between model outputs and human-written summaries. ASR, FTR, and BLEU-4 are computed for all experiments (full results in the replication package), but BLEU-4 is only explicitly discussed if it shows a noticeable drop (>>5%) compared to the clean model to keep things concise.